# A Systematic Review of NLP for Dementia- Tasks, Datasets and Opportunities

### v1

@@ Abstract
The close link between cognitive decline and language has fostered long-standing collaboration between the NLP and medical communities in dementia research. To examine this, we reviewed over 200 papers applying NLP to dementia-related efforts, drawing from medical, technological, and NLP-focused literature. We identify key research areas, including dementia detection, linguistic biomarker extraction, caregiver support, and patient assistance, showing that half of all papers focus solely on dementia detection using clinical data. However, many directions remain unexplored—artificially degraded language models, synthetic data, digital twins, and more. We highlight gaps and opportunities around trust, scientific rigor, applicability, and cross-community collaboration, and showcase the diverse datasets encountered throughout our review—recorded, written, structured, spontaneous, synthetic, clinical, social media-based, and more. This review aims to inspire more creative approaches to dementia research within the medical and NLP communities.
@@ 1 Introduction
There are multiple types of dementia, including Alzheimer’s (accounting for 60%-70% of cases111https://www.who.int/news-room/fact-sheets/detail/dementia), vascular dementia, Lewy-body dementia, and others. Although these differ in pathology—some manifesting as abnormal protein deposits and others as reduced blood flow to the brain—they often share common symptoms such as memory loss, confusion, behavioral changes, and language deterioration McKhann et al. (2011). In this review, we use the general term ’dementia’ to encompass all forms and levels of severity of the illness.
@@
Dementia is currently diagnosed through a battery of pathological tests (e.g., brain imaging, blood tests) and cognitive assessments, such as Naming and Verbal Fluency tests, which assess the ability to identify objects or generate words within specific categories Kaplan et al. (2001), and Picture Description tests, evaluating language and narrative skills through prompted descriptions of complex images Mueller et al. (2018). These cognitive assessments are typically conducted in face-to-face interviews with trained medical professionals, who look for linguistic markers of cognitive deterioration, such as repetitive speech, empty speech lacking meaningful content, disorganized narratives, word-finding difficulties, or a tendency to use indirect language and excessive descriptions (e.g., ’the thing you write with’ instead of ’pen’).
@@
Given the prominent role of language in our understanding and diagnosis of the disease, it is no surprise that there is substantial use of NLP methodologies in dementia research within both the NLP and medical communities. Cognitive assessments and similar diagnostics often result in recorded and transcribed data, offering a wealth of textual content that is well-suited for NLP analysis. This, combined with relatively new sources of data (such as social media posts or conversations with LLMs), enhance the potential of NLP to advance dementia research in various directions.
@@
Previous literature reviews on NLP for dementia focus on detection methods (Clarke et al., 2020; Saleem et al., 2022; Qi et al., 2023), with some specifically covering deep learning approaches to detection (Shi et al., 2023; Javeed et al., 2023). Other reviews extend beyond detection, addressing tasks like linguistic biomarker extraction (Gagliardi, 2024) or exploring the role of technology in enhancing patients’ life (D’Onofrio et al., 2017; Shahapure et al., 2022; Saragih et al., 2023; Peres and Campos, 2024), though these are not fully focused on NLP methods. A few reviews cover dementia-related data (De la Fuente Garcia et al., 2020; Yang et al., 2022) or, even more specifically, data from cognitive assessments like picture description tasks (Mueller et al., 2018).
@@
Our review stands out by covering the full spectrum of dementia-related NLP efforts, rather than focusing on specific aspects like data, tasks (e.g., detection), technologies (e.g., deep learning), data types (e.g., clinical), or communities (NLP vs. medical). Moreover, we created our review with NLP readers in mind, unlike prior reviews, which were primarily published in medical literature and aimed at clinicians (see Figure 3). Additionally, we review an extensive 205 papers, covering four scientific communities, categorized by publication venue: the medical community (e.g., Journal of Alzheimer’s Disease); the NLP community (e.g., ACL); the Speech community (e.g., InterSPEECH); and a broader technological community that is not necessarily NLP- or Speech-specific (e.g., Frontiers in Computer Science). In Section 2, we detail the construction of our cohort.
@@
In Section 3 we describe the main research verticals we encountered across our reviewed works, including linguistic biomarker extraction, caregiver support, patient assistance and, most prominently, dementia detection, comprising a striking 57% of the reviewed works focusing on this task alone. For each of these task families, we review motivations, current approaches, and potential future directions.
Our extensive review reveals a broader perspective on the datasets used for these tasks. We find that the vast majority of reviewed works—across all four task families—rely on a few well-known datasets. However, throughout our review we encountered many unique datasets of varying sizes, types, annotations, and purposes. Given the data-driven nature of NLP, we believe highlighting these datasets to the community could significantly advance research and open new avenues for practitioners. We discuss data gaps and opportunities in Section 4, and present a comprehensive table of the 17 datasets we encountered in Table 2.
Furthermore, in Section 4 we offer in-depth insights on the field’s scientific rigor and NLP’s potential to significantly shape medical research, provided trust between the two communities is established. We explore exciting research directions like personalized approaches, artificially degraded language models, and more.
@@
Our novel work primarily aims to inspire researchers from both fields to explore the vast potential of NLP in dementia research. We hope this paper provides a fresh perspective on the domain, emphasizing that the opportunities extend far beyond dementia detection. Whether developing clinical applications, analyzing disease progression, or alleviating caregiver burden, this review serves as a valuable resource for those looking to contribute to the fight against dementia.
@@ 2 Methodology
We searched for works applying NLP to dementia-related tasks across six data sources-ACL Anthology, PubMed, DBLP, IEEE Xplore, Springer, and Wiley. For a detailed description of our search queries along with our inclusion and exclusion criteria see Appendix A. In short, our focus was on peer-reviewed papers explicitly mentioning both dementia-related and NLP-focused terms in their titles or abstracts, and that are fully available in English. Many NLP studies on dementia use transcribed speech datasets with both audio and text; we exclude those focusing solely on voice signals without linguistic analysis.
Our screening resulted in 205 relevant papers, which we manually annotated to document their main contributions, datasets used, applied technologies (where applicable), and whether statistical significance was reported (see Appendix A).
We identified four distinct research verticals differing by task and motivation- dementia detection, linguistic biomarker extraction, caregiver support and patient assistance (see Section 3). Two additional, more general categories are literature reviews and dataset introduction papers, which we describe throughout our work. The number of works on NLP for dementia has been steadily growing (Figure 1), reflecting significant interest in the field. While we see a rise in works on patient assistance, caregiver support, and literature reviews, there is still an overwhelming focus on dementia detection compared to other tasks (Figure 2).
In the next section we dive into the four main task families we identify, discussing motivations, trends, and gaps. Naturally, the imbalance between the four tasks (particularly the fact that there are over 110 dementia detection papers while other task families have around 10-40 papers each) affects the depth of our analysis of specific dementia detection studies. Appendix B offers further details on the dementia detection works we reviewed.
@@ 3 Task Families
Can an algorithm accurately predict dementia from a provided text? Remarkably, 57% of the papers we reviewed address this exact question. This focus stems from two key factors. First, a tangible impact: NLP can significantly improve the diagnostic process patients undergo by making it faster, less invasive, and more affordable. Pathological changes can begin 15-20 years before cognitive symptoms are clearly noticeable to others Jack et al. (2010), so a quicker and more sensitive detection mechanism could enable earlier intervention. This may potentially slow disease progression and delay costly care, projected to reach $7 trillion in the U.S. by 2050.222https://www.prnewswire.com/news-releases/new-alzheimers-association-report-reveals-sharp-increases-in-alzheimers-prevalence-deaths-and-cost-of-care-300616178.html
@@
The second motivation to pursue dementia detection is a data-driven one. NLP research is inherently data-driven, and in this domain, data is often structured—spanning medical, audio, or textual modalities—annotated with categorical labels such as ’healthy’ or ’dementia.’ This structure perfectly suits classification algorithms, making dementia detection a well-defined challenge for researchers.
@@
NLP-based dementia detection is a research domain spanning over a decade, primarily relying on well-known datasets from the DementiaBank cohort Lanzi et al. (2023). Most used is the Pitt corpus Becker et al. (1994), offering transcribed recordings of cognitive assessments such as picture descriptions. The popular ADReSS and ADReSSo challenge datasets (Luz et al., 2021a, b, respectively) are derivatives of the Pitt corpus, offering refined transcripts and a demographically balanced sample. The CCC dataset Pope and Davis (2011) contains transcribed conversations on memories, daily life, and health. Most studies on detection utilize these transcribed conversations (or alternatively, use self-collected clinical notes), annotated with Dementia or Healthy Control labels.333Some datasets provide more granular annotations such as Dementia, Mild Cognitive Impairment, Alzheimer’s, etc.
@@
Across the 116 detection papers we reviewed, several algorithmic approaches emerged. Please note that we present representative citations here; for the complete list of references, see Appendix B. Before 2017, many studies relied on classic machine learning methods such as SVM and Random Forests using straightforward features like N-grams (Jarrold et al., 2014; Fraser et al., 2016a; Zhou et al., 2016; SANTOS et al., 2017), achieving accuracy rates around 85%. From 2018 onwards, there was a surge in transformer-based classifiers, particularly applied to the ADReSS and ADReSSo challenges published around that time (Pappagari et al., 2020; Edwards et al., 2020; Haulcy and Glass, 2021; Balagopalan et al., 2021). These methods pushed accuracy over 90%, representing the current state-of-the-art on the Pitt corpus and its subsets. Recently, large language models (LLMs) have been utilized for detection via prompt design, embedding extraction, and feature design (Agbavor and Liang, 2022; Wang et al., 2023a; Runde et al., 2024; BT and Chen, 2024; Bang et al., 2024).
@@
Given the extensive body of works and impressive classification results, one might assume that dementia detection is a solved task. However, (a) these results are mostly based on the Pitt corpus or its subsets—small, homogeneous datasets with very specific characteristics (structured conversation, English only, etc.); (b) only 24% of these works report statistical significance or assess the robustness of their findings; and (c) despite strong results, no NLP-based classification tool, to our knowledge, has advanced to a real-world application. We explore these gaps further in Section 4.
@@
Some researchers use NLP not for detection, but for various types of linguistic exploration. One motivation is to confirm, on a large scale, the existing knowledge we have about the language of cognitively impaired individuals (as reviewed by Clark et al., 2014; Fraser et al., 2016a; Voleti et al., 2019, and others). For instance, studies such as Orimaye et al. (2014), Rosas et al. (2019) and Ilias and Askounis (2022a) used NLP methods to solidify that linguistic behaviors associated with cognitive decline—such as repetitions, revisions, and simpler language—are much more common in cognitively impaired individuals than in healthy controls.
Other studies introduce novel linguistic metrics and markers to assess a speaker’s cognitive state. Examples include complexity scores from parse trees (Roark et al., 2007), idea density, which measures the efficiency of idea conveyance (Sirts et al., 2017), and the use of meta-semantic terms in picture descriptions, such as words implying emotion, emphasis, or opinion (Choi et al., 2019). The amount of disfluencies—including silent pauses, reformulations, and context switches (Adhikari et al., 2021; Farzana et al., 2022; Williams et al., 2023)—have also proven to be strong linguistic markers, even enabling the longitudinal tracking of disease progression (Martinc et al., 2021; Robin et al., 2023). These findings may influence NLP practices as they show that certain preprocessing steps, like removing repetitions and stop words, could erase valuable dementia-related signals. They may also impact the medical community by suggesting new metrics, like idea density, for novel cognitive assessments.
Notably, 70% of works in this task family report statistical significance—the highest ratio among all task families. This likely stems from the motivation to ensure that existing knowledge, or any new linguistic insights, are robust.
@@
With the growing number of dementia patients, the demand for caregivers rises. In the US alone, an estimated 11 million Americans serve as caregivers for those with dementia, contributing billions of hours of care each year.444https://www.alzint.org/about/dementia-facts-figures/ Research indicates that nearly half of these caregivers experience depression, and are at a higher risk of chronic health conditions.555https://www.caregiver.org/news/caregiver-isolation-and-loneliness/ NLP methods can help support these devoted caregivers by detecting emotional distress, providing answers to their concerns and offering companionship.
NLP for caregiver support, comprising less than 6% of our reviewed works, emerged after the COVID-19 outbreak, focusing on the emotional well-being of personal caregivers—i.e., family and friends- primarily through social media posts (Monfared et al., 2021). For example, Sunmoo et al. (2022, 2023) used graph-based topic modeling and sentiment analysis to show a shift in tweets posted throughout the pandemic—from practical care topics to emotional distress (e.g., depression, helplessness, elder abuse) and coping strategies (e.g., therapeutic reading). Posts beyond dementia-specific communities (Ni et al., 2022) and across social media profiles (Klein et al., 2022) reveal that caregivers express more than emotional distress, asking also for financial aid and legal advice.
Focusing on professional caregivers, Zhu et al. (2022b) analyzed clinical notes by nurses in aged care facilities. Disturbingly, they unveil dozens of distinct aggressive behaviors—including pushing, shouting, and using profane language—involving over 50% of dementia patients studied. Works like these highlight the potential of NLP in detecting when to assist caregivers, whether through advice, emotional support, or enhanced workforce training.
A growing body of research is exploring the potential of LLMs in providing caregivers with answers to dementia-related inquiries. Current studies show that models like ChatGPT offer relevant and factually correct advice and recommendations (Hristidis et al., 2023; Dosso et al., 2024), but often lack the depth of information available through Google Search or expert sources. Furthermore, LLMs fall short in offering the emotional support caregivers need when seeking to manage a loved one’s memory loss, confusion, or aggression (Aguirre et al., 2024). To bridge these gaps, Zaman et al. (2023) and Parmanto et al. (2024) explore the novel concept of LLMs fine-tuned specifically to address caregivers’ delicate emotional needs. These studies take a step toward AI-based caregiver companions, a field that is likely to gain traction with the rising popularity of LLMs.
@@
Individuals with dementia can live with the disease for up to 20 years,666https://www.alzheimers.org.uk/about-dementia/symptoms-and-diagnosis/how-dementia-progresses/later-stages-dementia and NLP solutions can significantly improve their daily lives. NLP can also help researchers decode the complex mental landscapes of patients, where agitation, depression, and cognitive decline intersect with daily experiences. Research indicates that depression affects 30-50% of dementia patients Lyketsos and Lee (2003), with 15% experiencing suicidal thoughts Naismith et al. (2022). This highlights the importance of works such as Fraser et al. (2016b) and Ehghaghi et al. (2022), who aim to detect depression among dementia patients. These studies reveal that relying solely on transcriptions of cognitive assessments makes it extremely difficult to detect depression among dementia patients, and that incorporating multi-modal data, particularly audio, enhances the results.
@@
Apart from assessing the mental well-being of patients, NLP can also aid in text simplification, making online information more accessible to those who often find it unreadable Espinosa-Zaragoza et al. (2023). Engineer et al. (2023) examined dementia-related texts online, revealing that the average readability level corresponds to 12 years of education—a level that many communities at higher risk for dementia may not have, given the research linking lower education levels to increased dementia susceptibility.777https://www.ncbi.nlm.nih.gov/pmc/articles/PMC6937498/ Additionally, the study reveals that the texts have a consistently negative tone, potentially affecting the mental state of patients who are already prone to depression. NLP tools are natural candidates to assist writers and content creators in producing more readable and supportive content.
@@
Researchers are also working to help patients sustain meaningful conversations and prevent communication breakdowns, shown to occur in over a third of interactions involving individuals with dementia, as shown by Chinaei et al. (2017). Both they and Green et al. (2012) have developed systems based on transcribed interviews to detect confusion and disfluent speech patterns, applying real-time repair strategies to reduce misunderstandings. These systems enhance the conversational experiences of dementia patients, improving their quality of life and reducing the burden on caregivers.
A growing body of research is exploring the potential of LLM-powered chatbots to improve social engagement, promote cognitive stimulation, and alleviate loneliness. Xygkou et al. (2024) and Qi (2023) explore GPT-based agents for reminiscence therapy (engaging in conversations about personal photos) as well as for general social interaction and accompaniment. Both studies found that participants valued the novelty and cognitive stimulation, reporting reduced loneliness and increased social support. However, participants noted limitations in handling emotionally sensitive conversations.
@@
LLMs for patients appear to be the next frontier in this task family, and many milestones have yet to be achieved—improved emotional capabilities, larger-scale experiments, and the development of evaluation benchmarks Gholizadeh et al. (2023). Additionally, as with all AI tools, building trust with patients and caregivers is essential. Gilman et al. (2024) proposes specialized GPT training for dementia patients as a way to help establish this trust. Treder et al. (2024) interviewed patients and caregivers about their views on such technologies, revealing concerns about bias, data privacy, and emotional intelligence, alongside a heartwarming, cautious optimism. As one patient experimenting with a chatbot noted: "I could talk to the robot longer than I could talk to a human… she didn’t tell me if I’ve repeated myself… she [the bot] didn’t think I was boring". Xygkou et al. (2024)
Summary

### v2

@@ Abstract
The close link between cognitive decline and language has fostered long-standing collaboration between the NLP and medical communities in dementia research. To examine this, we reviewed over 240 papers applying NLP to dementia-related efforts, drawing from medical, technological, and NLP-focused literature. We identify key research areas, including dementia detection, linguistic biomarker extraction, caregiver support, and patient assistance, showing that half of all papers focus solely on dementia detection using clinical data. Yet, many directions remain unexplored- artificially degraded language models, synthetic data, digital twins, and more. We highlight gaps and opportunities around trust, scientific rigor, applicability and cross-community collaboration. We raise ethical dilemmas in the field, and highlight the diverse datasets encountered throughout our review- recorded, written, structured, spontaneous, synthetic, clinical, social media-based, and more. This review aims to inspire more creative, impactful, and rigorous research on NLP for dementia.
@@ 1 Introduction
There are multiple types of dementia, including Alzheimer’s (accounting for 60%-70% of cases111https://www.who.int/news-room/fact-sheets/detail/dementia), vascular dementia, Lewy-body dementia, and others. Although these differ in pathology - some manifesting as abnormal protein deposits and others as reduced blood flow to the brain - they often share common symptoms such as memory loss, confusion, behavioral changes, and language deterioration McKhann et al. (2011). In this review, we use the general term ‘dementia’ to encompass all forms and levels of severity of the illness.
@@
Dementia is currently diagnosed through pathological tests (e.g., brain imaging, blood tests) and cognitive assessments, such as Naming and Verbal Fluency tests, which assess the ability to identify objects or generate words in specific categories Kaplan et al. (2001), and Picture Description tests, evaluating language and narrative skills through descriptions of complex images Mueller et al. (2018). In these face-to-face interviews, clinicians seek linguistic markers of cognitive decline, such as repetitive language, empty or disorganized speech, word-retrieval difficulties, and excessive descriptions ("the thing you write with" instead of "pen").
@@
Given the prominent role of language in our understanding and diagnosis of the disease, it is no surprise NLP methodologies are widely used in both the NLP and medical communities. Cognitive assessments and similar diagnostics often result in recorded and transcribed data, offering a wealth of textual content that is well-suited for NLP analysis. This, combined with relatively new sources of data (such as social media posts or conversations with LLMs), enhance the potential of NLP to advance dementia research in various directions.
@@
Previous literature reviews on NLP for dementia focus on detection methods (Clarke et al., 2020; Petti et al., 2020; Saleem et al., 2022; Parsapoor, 2023; Qi et al., 2023; Hiremath and Biradar, 2023; Vrindha et al., 2023), with some specifically covering deep learning approaches to detection (Shi et al., 2023; Javeed et al., 2023). Other reviews extend beyond detection, addressing tasks like linguistic biomarker extraction (Gagliardi, 2024) or exploring the role of technology in patients’ life (D’Onofrio et al., 2017; Shahapure et al., 2022; Saragih et al., 2023; Peres and Campos, 2024), though these are not fully focused on NLP methods. A few reviews are even more specific, covering dementia-related data (Mueller et al., 2018; De la Fuente Garcia et al., 2020; Yang et al., 2022).
@@
Our review stands out by covering the full spectrum of dementia-related NLP efforts, rather than focusing on specific aspects like tasks (e.g., detection), technologies (e.g., deep learning), data types (e.g., clinical), or communities (NLP vs. medical). Moreover, we created our review with NLP readers in mind, unlike prior reviews, which were primarily published in medical literature and aimed at clinicians (Figure 3). Additionally, our work is distinguished by its scope and coverage- we review 242 papers from four distinct scientific communities, categorized by publication venue: the medical community (e.g., Nature Medicine); the NLP community (e.g., ACL); the Speech community (e.g., Interspeech); and a broader technological community that is not necessarily NLP- or Speech-specific (e.g., Frontiers in Computer Science). In Section 2, we detail the construction of our cohort.
@@
In Section 3, we describe the main task families identified in our reviewed papers: linguistic biomarker extraction, caregiver support, patient assistance, and, most prominently, dementia detection, which accounts for over 56% of the cohort’s focus. For each task family, we review motivations, current approaches, and potential future directions.
Section 4 examines gaps and opportunities. We highlight, for instance, that the vast majority of studies - across all four task families - rely on a handful of well-known datasets, despite the existence of many unique datasets varying in size, type, and purpose (a summary of 17 dementia-specific datasets can be found in Table 2). We then dive into the field’s scientific rigor and explores NLP’s potential to influence medical research- provided trust between the two communities is established.
To provide readers with a concrete takeaway on future research directions, Section 5 outlines open challenges such as personalized LLMs for patients and caregivers, artificially degraded language models and many more. We conclude with the unique ethical considerations of the field (Section 6) and the limitations of our review (Section 7).
@@
Our work aims to inspire researchers from various fields to tap into the vast potential of NLP in dementia research. We hope to provide a fresh perspective on the domain, emphasizing that the opportunities extend far beyond detection. Whether developing clinical applications, analyzing disease progression, or alleviating caregiver burden, this review serves as a valuable resource for those looking to contribute to the fight against dementia.
@@ 2 Methodology
This review was conducted according to the PRISMA guidelines Moher et al. (2009). We searched for NLP- and dementia-related papers in titles, abstracts, and keywords across ACL Anthology, PubMed, DBLP, IEEE Xplore, Springer, and Wiley (full query details in Appendix A). An automatic screening ensured papers were full academic studies (rather than posters or theses), peer-reviewed, and in English. We then manually screened for eligibility, selecting studies that (1) focus on text as a primary modality, (2) use datasets that at least partially in English (to narrow our scope while still addressing multilingual studies), and (3) were relevant to this review. For instance, while Botros et al. (2020) mentions dementia and language models in its abstract and keywords, it focuses on smart home sensors, making it ineligible. Appendix A provides further examples of papers that do not meet our relevance criteria, alongside a fully detailed overview of our query, screening process, inclusion criteria, and a PRISMA flowchart.
Our screening resulted in 242 relevant papers, which we then manually annotated with their main contributions, datasets used, algorithmic methods and whether statistical significance was reported (where applicable). Among this cohort, we identified four distinct task families, described below.
@@ 3 Task Families
We identified four distinct research verticals differing by tasks and motivations- dementia detection, linguistic biomarker extraction, caregiver support and patient assistance (see Section 3). Two additional, more general categories are literature reviews and dataset introduction papers, which we describe throughout our work. The number of studies on NLP for dementia has been steadily growing (Figure 2), reflecting a significant interest in the field. While a rise in studies on patient assistance, caregiver support, and literature reviews, there is still an overwhelming focus on dementia detection (Figure 2). Naturally, this imbalance between the four tasks is reflected in our cohort - over 130 dementia detection papers versus 15-60 papers for each other task family - affecting the depth of our analysis of specific dementia detection studies. Therefore, Appendix B offers further details on the dementia detection papers we reviewed.
@@
Can an algorithm accurately predict dementia from a provided text? Remarkably, 56% of the papers we reviewed address this exact question. This focus stems from two key factors. First, a tangible impact: NLP can improve the diagnostic process by making it faster, less invasive, and more affordable. Pathological changes can begin 15-20 years before cognitive symptoms are noticeable to others Jack et al. (2010), and if NLP algorithms can detect early stages of dementia - e.g., mild cognitive impairment (MCI) - they may enable earlier intervention, potentially slowing disease progression and delaying costly care.
@@
The second motivation to pursue dementia detection is a data-driven one. NLP research is inherently empirical, and in this domain, data is often structured - spanning medical, audio, or textual modalities - annotated with categorical labels such as ‘healthy’ or ‘dementia.’ This structure perfectly suits classification algorithms, making dementia detection a well-defined challenge for researchers.
@@
NLP-based dementia detection has been an active research domain for over a decade, primarily leveraging well-known datasets such as the Pitt corpus Becker et al. (1994), part of the DementiaBank cohort Lanzi et al. (2023). The Pitt corpus provides transcribed recordings of cognitive assessments, such as picture descriptions. These transcribed text excerpts (e.g., ‘there’s a young boy, uh, going in a cookie jar, and there’s a lit… a girl…’) are annotated with the speaker’s cognitive state (e.g., Healthy, Mild Cognitive Impairment, Alzheimer’s) allowing for text-based classification. The transcriptions are also annotated with demographic information and, for some participants, other cognitive assessment scores. Two other widely used data sources, the ADReSS and ADReSSo challenge datasets (Luz et al., 2021a, b, respectively), are derivatives of the Pitt corpus, offering refined transcripts and a more demographically balanced sample.
Unlike these datasets, which build on structured conversations, the widely used CCC dataset Pope and Davis (2011) contains transcriptions of spontaneous conversations, about memories, health, and daily life (e.g., ‘hmm. This is my problem (…) maybe you could help me find it’). These four popular datasets (Pitt, ADReSS, ADReSSo and CCC) exemplify the most common data acquisition approach in dementia detection: transcribed conversations (whether from well-known datasets or self-collected clinical data) annotated with Dementia or Healthy Control labels.222Some datasets offer more granular annotations, such as Dementia, Mild Cognitive Impairment, Alzheimer’s, etc.
@@
Across the 136 detection papers we reviewed, several algorithmic approaches emerged. Please note that we present representative citations here; for the complete list of references, see Appendix B. Before 2017, many studies relied on classic machine learning methods like SVM and Random Forests using straightforward features like N-grams (Jarrold et al., 2014; Fraser et al., 2016a; Zhou et al., 2016; SANTOS et al., 2017), achieving accuracy rates around 85%. From 2018 onwards, there was a surge in transformer-based classifiers, particularly applied to the ADReSS and ADReSSo challenges published around that time (Pappagari et al., 2020; Edwards et al., 2020; Haulcy and Glass, 2021; Balagopalan et al., 2021). These methods pushed accuracy over 90%, which are currently state-of-the-art results on the Pitt corpus and its subsets. Recently, large language models (LLMs) have been utilized for detection via prompt design, embedding extraction, and feature design (Agbavor and Liang, 2022; Wang et al., 2023a; Runde et al., 2024; BT and Chen, 2024; Bang et al., 2024).
@@
Given the straightforward nature of the dementia detection (a text classification task), extensive body of research, and the impressive classification results, one might assume that dementia detection is a solved problem. From a modeling perspective, it has indeed evolved like other text classification tasks - from classic methods to word embeddings, neural networks, and LLMs. However, it remains fundamentally different, requiring tailored preprocessing, posing challenges in data collection from older adults, and demanding specialized linguistic metrics for evaluation - complexities we explore further in our review. As for the truly impressive results, they are (a) largely based on the Pitt corpus or its subsets - small, homogeneous datasets with specific characteristics (structured conversation, English only, etc.); (b) only 29% of studies report statistical significance or assess the robustness of their findings; and (c) despite strong performance, no NLP-based classification tool, to our knowledge, has been deployed in real-world applications. We explore these gaps further in Section 4.
@@
Some studies leverage NLP for linguistic exploration, rather than straightforward detection. They often wish to confirm, on a large scale, the existing knowledge we have about the language of cognitively impaired individuals. For instance, studies such as Orimaye et al. (2014), Rosas et al. (2019) and Ilias and Askounis (2022a) used NLP methods to verify that linguistic cues associated with cognitive decline - such as repetitions (e.g., ‘the… the…’), revisions (e.g., ‘the woman, uh, mother’), and overuse of pronouns (e.g. ‘she’ instead of ‘mother’) - are indeed behaviors significantly more common in texts produced by the cognitively impaired (Clark et al., 2014; Fraser et al., 2016a; Voleti et al., 2019).
Other studies introduce linguistic markers and metrics to assess a speaker’s cognitive state. For example, Roark et al. (2007) extracted complexity scores from parse trees, quantifying the extent to which sentences produced by cognitively impaired individuals are structurally and grammatically less complex. Sirts et al. (2017) calculated idea density to measure how efficiently dementia patients convey ideas, while Pompili et al. (2020a) quantified the number and order of topics mentioned by speakers in cognitive interviews. In other studies, such as Choi et al. (2019), the use of meta-semantic terms - words implying emotion, emphasis, or opinion - was analyzed in picture descriptions as an indicator of cognitive performance. The frequency of disfluencies, including silent pauses, reformulations, and context switches (Adhikari et al., 2021; Farzana et al., 2022; Williams et al., 2023), has also been identified as a strong linguistic marker, even enabling the longitudinal tracking of disease progression (Martinc et al., 2021; Robin et al., 2023).
Findings from this task family impact both the fields of NLP and medicine. From an NLP perspective, they underscore that the language of cognitively impaired individuals differs drastically from typical language, presenting unique challenges to standard NLP practices. For example, these studies show that certain preprocessing steps traditionally used in NLP, such as removing repetitions and stop words, could inadvertently erase dementia-related linguistic signals that are critical for tasks like dementia detection. From a medical perspective, these studies have the potential to challenge conventional scoring practices in cognitive assessments. One example is Prud’hommeaux et al. (2011), who explored narrative recall tasks - where participants are asked to retell a story after hearing it. The researchers found that scoring this cognitive assessment based on the presence of specific story elements may be more effective for detecting MCI than scoring based on the overall summary, which is the traditional approach, thus challenging the current perspective on this task.
Notably, 75% of papers in this task family report statistical significance- the highest ratio among all task families. This likely stems from the motivation to ensure that existing knowledge, or any new linguistic insights, are robust.
@@
With the growing number of dementia patients, the demand for caregivers rises. In the US alone, an estimated 11 million Americans serve as caregivers for those with dementia, contributing billions of hours of care each year.333https://www.alzint.org/about/dementia-facts-figures/ Research indicates that nearly half of these caregivers experience depression, and are at a higher risk of chronic health conditions Huang (2022). NLP methods can support these devoted caregivers by detecting emotional distress, providing answers to their concerns and offering companionship.
Around 6% of our cohort consists of studies on NLP for caregiver support, a field that emerged post-COVID-19. Early research focused on the emotional well-being of personal caregivers - family and friends - mainly through social media posts (Monfared et al., 2021; Azizi et al., 2024). For example, Sunmoo et al. (2022, 2023) used graph-based topic modeling and sentiment analysis to show that tweets posted throughout the pandemic shifted from practical care to emotional distress (e.g., depression, helplessness, elder abuse) and coping strategies (e.g., therapeutic reading). Posts beyond dementia-specific communities (Ni et al., 2022; Lal et al., 2023) and across social media profiles (Klein et al., 2022) reveal that caregivers express more than emotional distress, asking also for financial aid and legal advice.
Other studies focus on professional caregivers. One example is Zhu et al. (2022b), who analyzed clinical notes by nurses in aged care facilities. Disturbingly, their analysis shows dozens of distinct aggressive behaviors targeted towards the nursing team - including pushing, shouting, and using profane language - involving over 50% of dementia patients studied. This line of research highlights the potential of NLP in detecting when to assist caregivers, whether through advice, emotional support, or enhanced workforce training.
Another growing body of research is exploring whether LLMs could provide caregivers with expert answers to dementia-related inquiries. Current studies show that models like ChatGPT offer relevant and factually correct advice and recommendations (Hristidis et al., 2023; Dosso et al., 2024), but often lack the depth of information available through Google Search or expert sources. LLMs also fall short in offering the emotional support needed when a caregiver seeks to manage a loved one’s memory loss, confusion, or aggression (Aguirre et al., 2024). To bridge these gaps, Zaman et al. (2023) and Parmanto et al. (2024) explore the novel concept of fine-tuning LLMs to specifically address caregivers’ delicate emotional needs. These studies pave the way for AI-based caregiver companions, a field likely to gain traction with the growing popularity of LLMs. While groundbreaking, these agent-based solutions also present critical ethical dilemmas, explored in Section 6.
@@
Individuals with dementia can live with the disease for decades,444https://www.alzheimers.org.uk/about-dementia/symptoms-and-diagnosis/how-dementia-progresses/later-stages-dementia and NLP solutions can significantly improve their daily lives. NLP can also help researchers decode the mental landscapes of patients, where agitation, depression, and cognitive decline intersect. Research indicates that depression affects 30-50% of dementia patients Lyketsos and Lee (2003), with 15% experiencing suicidal thoughts Naismith et al. (2022). This highlights the importance of studies such as Fraser et al. (2016b) and Ehghaghi et al. (2022), who aim to detect depression among dementia patients. These studies demonstrate that it is extremely challenging to detect depression from the transcribed cognitive assessments alone, and that adding multi-modal data, particularly audio, drastically helps.
@@
Apart from assessing the mental well-being of patients, NLP can also aid in text simplification, making online information more accessible to those who often find it unreadable Espinosa-Zaragoza et al. (2023). Engineer et al. (2023) analyzed dementia-related texts online, finding an average readability level equivalent to 12 years of education - potentially inaccessible to many at-risk communities, given the link between lower education and dementia susceptibility.555https://www.ncbi.nlm.nih.gov/pmc/articles/PMC6937498/ The study also reveals a consistently negative tone in the texts (e.g., emphasizing the grim nature of the disease), potentially affecting the mental state of patients who are already prone to depression. NLP tools are natural candidates to help simplify texts, enhance readability, and rephrase to a more supportive tone. They can also assist writers and content creators ensure their work does not promote or reinforce stigma against patients Pilozzi and Huang (2020).
@@
Researchers are also trying to help patients sustain meaningful conversations and prevent communication breakdowns, shown to occur in over a third of interactions involving individuals with dementia. Studies such as Green et al. (2012) and Chinaei et al. (2017) have developed systems to detect confusion and disfluent speech patterns within transcribed interviews, suggesting real-time repair strategies to minimize misunderstandings. Such solutions may enhance conversations for dementia patients while also easing the burden on caregivers.
A growing body of research explores LLM-powered chatbots to enhance social engagement, provide cognitive stimulation, and reduce loneliness (Kostis et al., 2022; Xygkou et al., 2024; Qi, 2023; Gholizadeh et al., 2023). Studies show that participants appreciate the novelty and cognitive stimulation, reporting reduced loneliness and increased social support. However, they note some limitations such as handling emotionally sensitive conversations. Encouragingly, studies like Addlesee and Eshghi (2024) focus on improving chatbot patience and empathy, advancing the development of more suitable LLMs for patients.
@@
Additionally, as with all AI tools, building trust with patients and caregivers is essential. Gilman et al. (2024) proposes specialized GPT training for dementia patients as a way to help establish this trust. Studies such as Pacheco-Lorenzo et al. (2024) and Treder et al. (2024) interviewed patients, healthy individuals, and caregivers about their views on such technologies, revealing concerns about bias, data privacy, and emotional intelligence, alongside a heartwarming, cautious optimism. As one patient experimenting with a chatbot noted: "I could talk to the robot longer than I could talk to a human… she didn’t tell me if I’ve repeated myself… she [the bot] didn’t think I was boring". Xygkou et al. (2024)
Overview